OpenStudy (anonymous):

.Caroline's bedroom is the shape of a rectangular prism that measures 13 feet long, 10 feet wide, and 9 feet high. What is the surface area of her room

OpenStudy (anonymous):

how do i find the surface area?

jimthompson5910 (jim_thompson5910):

Area of Left Wall: A = LW A = 10*9 A = 90 Area of Right Wall = Area of Left Wall = 90 square feet -------------------------------------------------- Area of floor A = LW A = 13*10 A = 130 Area of ceiling = Area of floor = 130 square feet --------------------------------------------------- Area of front wall A = LW A = 13*9 A = 117 Area of back wall = Area of front wall = 117 square feet ====================================== Now add up the individual surface areas (Area of left wall) + ((Area of right wall) (Area of floor) + ((Area of ceiling) + (Area of front wall) + ((Area of back wall) ----------------------------------- 90 + 90 130+130 + 117+117 ----------- 180 260 + 234 ------------ 674 So the total surface area of the room is 674 square feet